The following data represents the number of packages of 5 popular brands of cigarettes sold by a supermarket on 8 randomly selected days:

 

Brand

A

B

C

D

E

21

35

32

28

14

47

25

38

35

12

27

41

19

23

31

20

45

60

33

36

31

40

43

48

32

53

29

42

40

23

35

42

45

29

31

22

36

29

42

30

 

 

Perform an analysis of variance, at 0.05 level of significance, and determine whether or not the 5 brands sell, on the average, the same number of cigarettes at this supermarket. (15 points)

 

 



1
Expert's answer
2021-12-21T16:09:30-0500


One Way ANOVA

"H_0: \mu_1=\mu_2=\mu_3=\mu_4=\mu_5" , 5 brands sell, on the average, the same number of cigarettes at the supermarket

"H_a: not\ (\mu_1=\mu_2=\mu_3=\mu_4=\mu_5)" , 5 brands sell, on the average, not the same number of cigarettes at the supermarket


F statistic:

F = MSG / MSE = 2.725

where:

Mean Square between groups:

MSG = SSG / (k - 1) =237.24

k = 5 is number of groups

Sum of Squares between groups:

"SSG=\sum n_i (\overline{x}_i-\overline{x})^2=948.98"

Mean Square within groups:

MSE = SSE / (n - k) = 87.05

n = 45 overall sample size

Sum of Squares within groups:

"SSE=\sum (n_i-1)s^2_i=3482"


Degrees of Freedom:

between groups: "k-1=4"

within groups: "n-k=40"

critical value:

"F{crit}=2.606"


Since "F>F{crit}" we reject null hypothesis. 5 brands sell, on the average, not the same number of cigarettes at the supermarket.